What is the rms value of current i=3+4sin(wt+pi\3)? Please give step by step solution.

Grade:12

1 Answers

prateek
7 years ago
i=Asin(wt+fi)
A=peak current
comparing from your question A=4 
I(r m s)=A/21/2 
I(r m s)=23/2
i hope my answer is correct . I hope that u understood. Please contact for any other queries.
THANKS AND STUDY HARD!!!!!